OpenStudy (anonymous):

Have a question: -x + 3y = 12 and y = 5 Would x = 3 or -3?

jimthompson5910 (jim_thompson5910):

y = 5 so -x + 3y = 12 turns into -x + 3*5 = 12 after you plug in y = 5

jimthompson5910 (jim_thompson5910):

-x + 3*5 = 12 -x + 15 = 12 -x = 12 - 15 -x = -3 We want x, not -x, so change the sign of both sides to get x = 3

OpenStudy (anonymous):

Thank you Jim, I thought it was 3 but I wasn't completely sure, I forgot to change -x to a positive. Thanks again!

jimthompson5910 (jim_thompson5910):

you're welcome, glad to be of help